### v1

# DRAGON: Dynamic RAG Benchmark On News

Abstract
Retrieval-Augmented Generation (RAG) is a widely adopted approach for improving the factuality of large language models (LLMs) by incorporating external knowledge at inference time. Although there exist multiple RAG benchmarks for English, evaluation resources for other languages, including Russian, remain scarce and static, failing to capture the dynamic nature of real-world deployments.
In this work, we present DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating RAG systems in Russian on a changing news corpora. DRAGON is built upon a regularly updated corpus of Russian news and public documents and supports comprehensive evaluation of both the retriever and generator components. Question generation is performed automatically with the use of Knowledge Graph constructed from the corpus and enables the extraction of four core question types aligned with distinct subgraph patterns. We release a complete evaluation framework comprising the pipeline for automatic question generation, evaluation scripts, which are potentially reusable for other languages and multilingual settings, and benchmark data. We also launch a public leaderboard to encourage community participation and comparison.
DRAGON: Dynamic RAG Benchmark On News
1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful instrument for enhancing the domain adaptation and factuality of large language models (LLMs) by incorporating external knowledge retrieved at inference time. This approach enables more up-to-date and grounded responses without the need for costly re-training. As RAG-based systems expand to applications such as open-domain QA, customer support, and enterprise search, their standardized evaluation becomes crucial for assessing reliability and utility.
While several RAG benchmarks have been proposed for the English language, there are significantly fewer evaluation resources for other languages, including Russian.
In this work, we focus on the evaluation of RAG systems for the Russian language. We introduce DRAGON: Dynamic RAG Benchmark On News (Fig. 1) a novel benchmark that reflects realistic usage patterns by leveraging a regularly updated knowledge base derived from current news sources. To foster transparency and community engagement, we publicly release an evaluation framework which comprises the codebase for automatic question generation, evaluation scripts, and a dynamic leaderboard to track progress on RAG-based systems in Russian. While the benchmark targets Russian, the framework is potentially extendable to other languages and multilingual scenarios, making it broadly applicable. Our contributions are as follows:
(i) We propose DRAGON111The video demonstration of the evaluation tool is available on YouTube., the first RAG benchmark with a regularly updated knowledge base based on news corpora for Russian, designed to evaluate RAG systems in a dynamic setup.
(ii) We release an open-source evaluation framework222The framework is released under the MIT license: https://github.com/RussianNLP/DRAGON comprising a reusable question generation pipeline and evaluation scripts, enabling reproducible experimentation and easy integration of new models and retrieval components. By design, it can potentially be adapted to other languages and multilingual settings, broadening its applicability beyond Russian.
(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
2 Related Work
Proper assessment of RAG systems requires holistic benchmarks that evaluate both the retrieval and the generation components. For English, a growing body of research has focused on designing such benchmarks across a range of reasoning and grounding challenges.
KILT Petroni et al. (2021) provides a unified benchmark over a fixed Wikipedia snapshot, with an emphasis on source attribution and retrieval grounding. Conversational settings with multi-turn dialogue scenarios are investigated in mtRAG Katsis et al. (2025) and RAD-Bench Kuo et al. (2025).
The factual consistency of generated responses in RAG has become a critical challenge. CRAG Yang et al. (2024) introduces a comprehensive benchmark based on five main RAG characteristics. RAGAS Es et al. (2024) proposes a reference-free evaluation suite that captures context relevance, faithfulness, and answer completeness, along with an open-source API for practical benchmarking.
Despite these advances, most of the RAG benchmarks focus on English. Evaluation for other languages remains limited. NoMIRACL Thakur et al. (2024) provides multilingual QA tasks in 18 languages and analyzes retrieval failures and hallucination rates. For Russian, RusBEIR Kovalev et al. (2025) offers retrieval-only evaluation, but a complete end-to-end RAG benchmark has yet to be developed.
To address this need we present DRAGON – a dynamic, regularly updated benchmark for Russian-language RAG systems based on real-world, shifting corpora.
3 System Demo
The benchmark is designed to evaluate RAG systems in a dynamically evolving news domain. Its architecture prioritizes modularity, automation, and reproducibility while addressing the core challenges Yu et al. (2024) in the RAG evaluation landscape such as temporal aspects of information, dynamic and vast knowledge sources, and factuality and faithfulness in generation.
Architecture of the Dynamic Benchmark
The entire pipeline of the benchmark architecture is shown in Fig. 2. Below, each step is described in more detail.
Data Acquisition and Processing: We have maintained a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx. A for the news sources list) on a daily basis. Newly parsed content is synchronized into our storage. To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
This design ensures the benchmark reflects evolving real-world distributions and mitigates the risks of overfitting to static datasets. The pipeline further ensures that newly surfaced topics and entities from the news stream are constantly incorporated into the benchmark.
QA Dataset Formation: The process of creating the questions and answers based on the updated increment of news data is described in detail in Sec. 4. The pipeline transforms generated QA pairs into several HF datasets, which form the core of our benchmark:
Public Texts: Contains cleaned source documents. Each item is assigned a public_id to enable matching without exposing the true internal IDs.
Public Questions: Contains only questions, indexed via public_id to obfuscate alignment and encourage retrieval.
Private Texts Mapping: Used only for evaluation purposes. It contains internal ids and the corresponding public_ids to enable accurate mapping during metric computation.
Private QA: Provides canonical ground-truth answers for generative evaluation.
In addition to these main datasets, we provide a separate set of Sandbox Datasets with the exact same structure as the main ones. All four sandbox datasets are fully public. Their purpose is twofold: (1) to transparently demonstrate the full structure and intended usage of the benchmark, and (2) to allow users to validate their RAG systems locally without submitting results to the validation portal.
These sandbox datasets can be evaluated using the rag_bench client library, which supports the same retrieval and generative metrics as those used by the official validation portal (except for judgment-based metrics). This enables convenient local experimentation, debugging, and reproducibility.
All datasets are versioned and uploaded to Hugging Face with incrementally updated revision numbers. This versioning mechanism ensures reproducibility and provides users with stable snapshots for further experimentation.
User experience
To facilitate seamless evaluation for users, we provide a PyPi-hosted Python library rag_bench, which offers an interface to:
Fetch the latest version of the public datasets by dynamically resolving the latest Hugging Face revision; Observe the RAG system baseline which can be adopted for the target one; Evaluate RAG system and package results for submission; Submit results via API to our evaluation portal; Calculate retrieval and generative metrics locally using the sandbox datasets.
User workflow includes loading public data, applying a custom RAG pipeline, and collecting results in the following form:
"0": {
"found_ids": [17, 69, 69, 22, ...],
"model_answer": "Answer": "Top"
These results encode both the retrieved public_ids and the generated answers, decoupling the user’s model output from any private evaluation artifacts. This separation allows secure evaluation without exposing ground-truth data.
Validation Portal
Submitted results then are sent to the Validation Portal — a Flask-based backend with a Single Page Application written in Vue as a frontend that performs secure evaluation using the private datasets. The portal evaluates submissions using private datasets and prepares evaluation results for admin approval before publishing. Importantly, users submit only their results — all ground-truth data remains internal.
Leaderboard and Auto-Evaluation
A Hugging Face Gradio Space serves as the public Leaderboard (see Appx. B for a screenshot). Results are committed in a version-controlled results.json file, automatically updated by the validation portal upon approval.
To reduce latency and improve benchmarking coverage, we support automatic evaluation for selected pre-approved baselines, which include several popular LLMs and retrieval embedding models (see Appx. C for the details). The results are computed via the same rag_bench client.
3.1 Versioning Strategy
Given the dynamic nature of the benchmark, versioning plays a critical role in ensuring meaningful comparisons. Each evaluation result is tied to a specific dataset revision. On the leaderboard, users can view results for a single dataset version or toggle an “Actual Versions” mode to aggregate results across recent revisions.
Dataset versioning is performed automatically based on the last available version on Hugging Face. The version number follows a semantic format, e.g., 1.10.0. For each new release, the middle segment of the version is incremented, resulting in a new version such as 1.11.0, which is then uploaded to Hugging Face. This approach ensures consistent, chronological dataset updates while preserving backward compatibility for previously published results.
Note that sandbox datasets are not updated on a regular basis. They serve as a static reference set for demonstration and local validation purposes.
4 Methodology
The Data Generation pipeline (Fig. 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation. The KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph. The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
Knowledge Graph Extraction
To enable greater control over question generation, we developed a Knowledge Graph Extraction module, inspired by Chepurova et al. (2024).
At the first stage, the LLaMa 3.3 70B Instruct444https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Grattafiori et al. (2024) is used for the extraction of triplet-candidates. Each triplet consists of a head entity or subject, a relation, and a tail entity or object.
All the extracted entities are used to match with entities from Wikidata555We use Russian subgraph of Wikidata as a base one. Vrandečić and Krötzsch (2014). To perform this, we search by API in this open knowledge base for the entities extracted in the previous step. Afterward, the found entities are vectorized using an embedding model and saved as a vector database. Then, for each entity extracted in the first step, we list five of the closest entity candidates in the vector representation base.
At the final stage, we use the same LLaMa 3.3 70B Instruct for normalization and unification of the extracted entities and relations. For each entity and relation in the triplet, the language model analyzes the list of candidates obtained at the previous stage and decides on the possibility of replacing the original entity or relation with one of the candidates. The triplet context is taken into account by adding an original text. This approach allows the unification of naming variants for the same entity, which is critical for the subsequent construction of the knowledge graph.
We aim to create a graph that contains novel information. To ensure this, we discard any triplets that are completely matched with saved entities and relations from the open knowledge base (Wikidata). We assume that facts not included in Wikidata are novel and thus unlikely to have been seen by the language model.
Question Generation
The question generation stage begins with subgraph extraction. We systematically sample all subgraphs that match one of the four predefined structural templates, each corresponding to one of the question types. We employ a typology of four question types, inspired by Yang et al. (2024): simple, set, multi-hop, and conditional questions (see Appx. D for a detailed description).
A subgraph is passed to the language model, which is instructed to generate both a natural language question and a corresponding answer. The question is formulated as a coherent sentence, while the answer consists of one or more entities derived from the subgraph.
QA Filtering
The dataset requires filtering to improve the quality of question-answer pairs and mitigate language model errors. Firstly, the linguistic acceptability of each question is measured using a model trained on the RuCoLa dataset666https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola Mikhailov et al. (2022) model, which scores sentences based on their grammatical correctness and fluency.
Secondly, we extract Named Entities from the original document, solving the NER task with Natasha library. We look for these entities in generated question-answer pairs. The pairs that do not contain named entities are discarded to avoid examples relying on general knowledge. The questions are passed to the small LLMs (namely, we use Qwen 2.5 7B Instruct Team (2024), Llama 3 8B Grattafiori et al. (2024)) without context to detect the simplistic pairs that do not challenge the language model.
Thirdly, each generated pair is checked for compliance with the source subgraph. Either question or answer should contain an entity from the subgraph. To ensure this, we search the closest by Levenshtein distance Levenshtein et al. (1966) entity from the subgraph. We consider an entity to be found if the smallest distance from the entities in the subgraph is under the predefined threshold. We assume that the generated set should contain all the entities from the subgraph and not contain any extra ones. Otherwise, there is a generation error or the question type mismatch (see Appx. E).
The final stage uses the LLM-as-Judge approach to measure the question and answer quality. The pretrained judge POLLUX 7B Martynov et al. (2025) specifically trained for a fine-grained evaluation in Russian is adopted for this task. The model evaluates each sample against 8 generative criteria developed to assess the quality of the question-answer pair and its context dependence (the set description is given in Appx. F). As a result, questions of low quality are excluded from the final test set. The evaluation process is described in Appx. G. After the filtering stage 150 questions for each category form the final test dataset.
5 Experimental Setup
To construct our experimental RAG systems, we used the LangChain framework777https://pypi.org/project/langchain/. All texts from the Public Texts dataset are split into chunks of length 500 with an overlap of 100 characters. Each chunk is vectorized using the retrieval model of the evaluating RAG system with the corresponding document prefixes (see Appx. H.1 for their exact formulation), and the resulting vectors are stored in a vector database.
During the search phase, we use the prompted retrieval model to find five of the most relevant texts that match the user’s query. Retrieved chunks are incorporated into a prompt (see Appx. H.2) provided to the LLM of the evaluated RAG system. If the total length of the filled-in prompt exceeds the model’s maximum context length, the contextual information is truncated to the required size. To accelerate LLM inference, we utilize the vLLM framework888https://github.com/vllm-project/vllm Kwon et al. (2023). Other configuration details are described in Appx. H.3.
Scoring metrics
The performance of retrieval is measured with the 3 main metrics: Hit Rate, Recall, and NDCG. End-to-end RAG-system evaluation is performed via ROUGE-L, Substring Matching (SM), and Judge Score. See Appx. H.4 for their description.
6 Experiments
Question Quality Evaluation
To assess the quality of the generated question-answer pairs, a human evaluation study is conducted. Each QA pair from Sandbox Datasets (Sec. 2) is independently evaluated by 3 expert annotators along the evaluation criteria from Appx. F. Annotators were asked to mark each pair as “Good” or “Not Good” with respect to each criterion (an example of the assessment instruction is described in Appx. I). To account for potential subjectivity in judgment, we considered a QA pair to be acceptable with the majority vote.
Tab. 1 shows the proportion of QA pairs considered good for each dataset version and each evaluation criterion. The results establish the high quality of generated questions and significant context dependency. The answer evaluation proved the prevalence of correct answers, while the answer uniqueness is lower, so the ground truth answer can be substituted with another entity from the text. This fact exhibits the importance of LLM-as-Judge evaluation for RAG systems to avoid rephrasing influence.
| Criterion | Apr | May | Jun | |
| Question | Question Literacy | 0.96 | 0.97 | 0.99 |
| Clarity | 0.99 | 1.00 | 1.00 | |
| Naturalness | 0.98 | 0.96 | 0.97 | |
| Context Sufficiency | 0.98 | 0.98 | 0.99 | |
| Context Necessity | 0.95 | 0.97 | 0.98 | |
| Answer | Correctness | 0.95 | 0.92 | 0.96 |
| Uniqueness | 0.76 | 0.78 | 0.80 | |
| Answer Literacy | 0.79 | 0.71 | 0.75 |
Retrieval Evaluation
Retrieval evaluation results presented in Tab. 2 demonstrate consistently strong performance across all evaluated retriever models. Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par except for MRR, where E5 Mistral7b Instruct beats its competitor. However, mE5lInstruct and FRIDA also perform competitively.
| Retriever | Hit Rate | Recall | NDCG |
| FRIDA | 0.88 | 0.84 | 0.85 |
| mE5 | 0.90 | 0.87 | 0.86 |
| Qwen 3 | 0.93 | 0.90 | 0.89 |
| E5 Mistral | 0.93 | 0.89 | 0.87 |
End-to-End System Evaluation
Tab. 4 presents general results of the full RAG-system evaluation and detailed criteria-wise evaluation scores are given in Appx. J.
First, it can be seen that the choice of the retrieval model plays a crucial role and same as in Sec. 6. Qwen3 and E5 Mistral show the strongest results. Second, it should be noted that the general LLM ranking remains the same with every retrieval, with Qwen 2.5 and Zero Mistral heading the list by Judge Score and SM, and Gemma 3 over-performing other competitors by Rouge-L.
Overall, the results show that classic metrics such as Rouge-L are not objective enough and do not allow evaluating all aspects of the RAG task. In general, system scores positively characterize DRAGON as being complex enough for modern RAG-systems, allowing researchers to evaluate their capabilities at a high level. In the future, we also plan to complexify Judge Evaluation criteria, thus providing an opportunity for an adequate assessment of more advanced models than those that exist nowadays and avoiding the danger of the benchmark being solved.
| Retriever | LLM | Rouge-L | SM | Judge Score |
| FRIDA | Qwen 2.5 | 0.42 | 0.47 | 0.83 |
| Qwen 2.5 | 0.33 | 0.43 | 0.80 | |
| Ruadapt Qwen | 0.40 | 0.45 | 0.82 | |
| Zero Mistral | 0.44 | 0.45 | 0.83 | |
| Gemma 3 | 0.44 | 0.44 | 0.81 | |
| Gemma 3 | 0.40 | 0.47 | 0.81 | |
| Qwen 3 | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
| Qwen 2.5 | 0.36 | 0.45 | 0.83 | |
| Ruadapt Qwen | 0.42 | 0.48 | 0.86 | |
| Zero Mistral | 0.45 | 0.48 | 0.86 | |
| Gemma 3 | 0.47 | 0.47 | 0.84 | |
| Gemma 3 | 0.43 | 0.49 | 0.85 | |
| E5 Mistral | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
| Qwen 2.5 | 0.36 | 0.46 | 0.82 | |
| Ruadapt Qwen | 0.41 | 0.47 | 0.85 | |
| Zero Mistral | 0.45 | 0.47 | 0.85 | |
| Gemma 3 | 0.46 | 0.47 | 0.83 | |
| Gemma 3 | 0.41 | 0.48 | 0.85 | |
| mE5 | Qwen 2.5 | 0.43 | 0.49 | 0.85 |
| Qwen 2.5 | 0.34 | 0.44 | 0.82 | |
| Ruadapt Qwen | 0.39 | 0.46 | 0.83 | |
| Zero Mistral | 0.44 | 0.46 | 0.84 | |
| Gemma 3 | 0.46 | 0.46 | 0.83 | |
| Gemma 3 | 0.41 | 0.48 | 0.83 |
7 Conclusion
We presented DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating retrieval-augmented generation systems in Russian. DRAGON reflects real-world deployment settings by leveraging a regularly updated knowledge base and focuses on the recurrent evaluation of both retriever and generator components. Our benchmark addresses the current lack of standardized RAG evaluation tools for the Russian language. We release the benchmark, the framework comprising a question generation pipeline and evaluation scripts, and launch a public leaderboard, to support reproducible, transparent, and community-driven research. In the future, with the evolving capabilities of RAG systems, we plan to extend the benchmark by introducing new question types, refining the LLM-as-Judge criteria. In addition, we aim to open-source previous snapshots of the evolving datasets to support reproducibility and foster further community research.
We hope DRAGON will serve as a foundation for future work on multilingual and dynamic RAG systems.
Limitations
While the proposed benchmark provides a valuable framework for evaluating retrieval-augmented generation (RAG) systems, several limitations should be acknowledged:
Source Diversity
The benchmark primarily relies on the available documents from a specific domain (news), which may not fully capture the diversity of real-world information retrieval and generation tasks. Expanding the dataset range could enhance the benchmark’s applicability across different domains.
Language Diversity
The proposed benchmark consists entirely of Russian language documents and questions. Although the methodology itself could be easily applied to any other language, in the current state the only one language is presented.
Evaluation Metrics
The chosen evaluation metrics, such as ROUGE, which is essentially an n-gram precision, predominantly focus on surface-level matching. These metrics may not adequately reflect the semantic and pragmatic aspects of the generated content, and have limited correlation with human judgment Deutsch et al. (2022). LLM as Judge evaluation is designed to mitigate a semantic gap of n-gram based metrics. However, RAG benchmark requires specific criteria to catch system performance details. Building more adapted judge models can improve the quality of the assessment.
Domain-Specific Challenges
RAG systems might perform differently across various domains due to domain-specific complexities and knowledge structures. The benchmark does not currently address these nuances, which could hinder its ability to generalize across distinct fields like medicine, law, or general knowledge.
Retriever-Generator Synergy
The interactions between retrieval and generation components are complex and dynamic. Our benchmark does not deeply explore how different configurations and synergistic interactions affect performance, possibly oversimplifying nuances that can significantly impact results.
Human Evaluation
The benchmark primarily relies on automated metrics, which may not align perfectly with human judgments of quality and relevance. While we acknowledge the role of human evaluation, it was not feasible to incorporate it extensively into this iteration of the benchmark.
Scalability and Efficiency
The computational resources required for comprehensive testing can be substantial, potentially restricting the accessibility of the benchmark to groups with extensive computational infrastructure.
Rapid Technological Advancements
The field of RAG systems is rapidly evolving, with new models and techniques emerging frequently. The benchmark may quickly become outdated unless regularly updated to incorporate recent advancements and methodologies.
Addressing these limitations in future work could involve developing more comprehensive, diverse datasets, incorporating a broader range of evaluation metrics, and continuously adapting the benchmark to reflect the state-of-the-art in RAG systems. Additionally, exploring detailed interactions between retrieval and generation components and integrating more human evaluation into the assessment process could provide deeper insights and improve the robustness of the benchmark.
Ethical consideration
In developing and utilizing the retrieval-augmented generation (RAG) systems benchmark, several ethical considerations have been taken into account to ensure responsible and fair use of the technology:
Bias and Fairness
Given that RAG systems are influenced by the data they are trained and tested on, it’s crucial to address the potential for bias in retrieval and generation processes. Our benchmark highlights these concerns by incorporating evaluation metrics that identify and measure biases in model outputs. Future iterations aim to include datasets specifically designed to stress-test and mitigate bias.
Data Privacy
The use of real-world datasets in RAG systems poses privacy risks, particularly concerning personally identifiable information (PII). We ensure that datasets included in the benchmark are sourced following strict privacy regulations and guidelines, and we encourage the anonymization of any PII to safeguard user privacy.
Content Quality and Misinformation
RAG systems can potentially generate or propagate misinformation if not properly managed. Our benchmark assesses models on their ability to produce accurate and reliable content, and we emphasize the importance of retrieval sources that are reputable and verifiable to minimize risks associated with misinformation.
Transparency and Explainability
Understanding the decision-making process of RAG systems is critical for trust and accountability. The benchmark encourages the development of models that offer insights into their retrieval and generation processes, promoting transparency and explainability.
Unintended Consequences
The application of RAG systems can have unintended societal impacts, such as fostering dependency on AI for decision-making or influencing cultural narratives. Researchers and developers are encouraged to consider these broader implications and involve interdisciplinary perspectives in assessing the impact of their systems.
Access and Inequality
High computational demands of RAG systems can exacerbate the divide between well-resourced organizations and smaller entities or individuals. Our benchmark advocates for the creation of more efficient models that democratize access and enable wider participation in developing and utilizing RAG technology.
Responsible Usage
Educating users and stakeholders about the capabilities and limitations of RAG systems is vital to prevent misuse. Our research promotes guidelines and best practices to ensure that these technologies are used responsibly and ethically.
By acknowledging and addressing these ethical considerations, our aim is to contribute positively to the development and deployment of retrieval-augmented generation systems, ensuring they serve society in a beneficial and responsible manner. Future work will continue to refine these frameworks to address emerging ethical challenges as the field evolves.
AI-assistants Help
We improve and proofread the text of this article using Writefull assistant integrated in Overleaf (Writefull’s/Open AI GPT models) and GPT-4o999https://chatgpt.com, Grammarly101010https://app.grammarly.com/ to correct grammatical, spelling, and style errors and paraphrase sentences. We underline that these tools are used strictly to enhance the quality of English writing, in full compliance with the ACL policies on responsible use of AI writing assistance. Nevertheless, some segments of our publication can be potentially detected as AI-generated, AI-edited, or human-AI-generated.
Appendix A News Data Sources
For dataset formation, we rely on content from several well-established Russian news websites111111All data is used in full compliance with legal requirements and ethical standards, under a formal agreement with Rambler. The collection process ensures respectful use of content without infringing on the rights of publishers or individuals.:
Appendix B Leaderboard Overview
Fig. 4 shows an overview of the leaderboard.
Appendix C Baseline Details
We evaluate open-source LLMs within 70B size121212The size limit is introduced to ensure the feasibility of multi-model evaluation under the compute budgets. which score best on the MERA benchmark131313https://mera.a-ai.ru/en/text/leaderboard, valid for July 1, 2025. Fenogenova et al. (2024) (see Tab. 4 for their description) and several popular embedding models which show strong results on the retrieval task on ruMTEB or Multilingual MTEB141414https://huggingface.co/spaces/mteb/leaderboard valid for July 1, 2025. Snegirev et al. (2024); Enevoldsen et al. (2025).
| Model | Size | Hugging Face Hub link | Citation |
| Qwen 2.5 | 32B | Qwen/Qwen2.5-32B-Instruct | Team (2024) |
| Qwen 2.5 | 32B | Qwen/Qwen2.5-32B-Instruct | Team (2024) |
| Ruadapt Qwen | 32B | msu-rcc-lair/RuadaptQwen-32B-instruct | Tikhomirov and Chernyshev (2024) |
| Zero Mistral | 24B | ZeroAgency/Zero Mistral-24B | – |
| Gemma 3 | 12B | google/gemma-3-12b-it | Team (2025) |
| Gemma 3 | 27B | google/gemma-3-27b-it | Team (2025) |
| Model | Size | Hugging Face Hub link | Citation |
| FRIDA | 823M | ai-forever/FRIDA | – |
| Qwen 3 | 8B | Qwen/Qwen3-Embedding-8B | Zhang et al. (2025) |
| E5 Mistral | 7B | intfloat/e5-mistral-7b-instruct | Wang et al. (2023) |
| mE5 | 560M | intfloat/multilingual-e5-large-instruct | Wang et al. (2024) |
Appendix D Question Types
We use the typology of four question types inspired by Yang et al. (2024). Below each type is described in more detail:
Simple
These type correspond the most simple questions based a single fact mentioned in one or several texts. They are based only on one relation from the graph: the predicate and one of the entities involved in the relation are used to compose the question, and the second entity becomes the answer.
relations: (Morty Smith | voice | Keisuke Chiba)
question: Who voiced Morty Smith?
answer: Keisuke Chiba
Set
Set questions test the RAG system’s ability to align information from several texts. They are based on a one-to-many subgraphs in which the number of triplets share relation and either object or subject. The question is generated using shared entity and relation. The answer consists of all other entities in the subgraph.
relations: (Ryan Otter | composed music for | Method), (Ryan Otter | composed music for | Trigger)
question: What projects has Ryan Otter composed music for?
answer: Trigger, Method
Multi-Hop
Multi-hop questions evaluate the system’s ability to reason in a multistage manner. The corresponding subgraph is a pair of triplets, intersecting at a single entity. The question is constructed similarly to a simple question; however, the repeated entity must not be mentioned in the question. It is used as a bridge-entity, which is described in question as a reference extracted from another triplet.
relations: (FAW | country of origin | China), (FAW | number of cars sold in 2023 | 2139)
question: In which country is the company located that sold 2139 cars in 2023?
answer: China
Conditional
Conditional questions are the extension of multi-hop questions with the same underlying subgraph of a pair of triplets, intersecting at a single entity. However, for a conditional question, both facts are used to form the question, while the repeated entity becomes an answer.
relations: (Roman Miroshnichenko | performed at | M-bar), (Roman Miroshnichenko | met with | Dmitry Dibrov)
question: Who performed at M-bar and met with Dmitry Dibrov?
answer: Roman Miroshnichenko
Appendix E Graph correspondence filtering
The question-answer pairs were compared to the graph used for generation. Each node in the graph (entity) is assigned 2 coefficients: question presence and answer presence. It is evaluated as the scaled Levenshtein distance between the name of the entity and the closest substring from the question and answer. These values allow us to check that all entities have been mentioned correctly.
In the graphs for Set and Conditional question types, the positions of every entity are strictly determined. The algorithm averages the presence coefficients of entities implied to be in the same part of the output. If any of these values is lower than the threshold, it indicates incorrect generation. For Simple questions each entity can appear in both parts of the output, although the entity must be mentioned once in the question-answer pair. The presence coefficients were averaged over all entities from the subgraph, then 5% highest and lowest values were filtered out. Multi-Hop questions inherit the same process for nodes having only one connection in the subgraph. The bridge entity that has two connections should not be mentioned in the model output. A high value for any of the presence coefficients for this entity demonstrates a question-type violation.
Appendix F Question-Answer Evaluation Criteria
This section describes the question-answer evaluation criteria used on the final question filtering stage. These criteria were developed to assess the general quality and naturalness of the question, its context dependence, and the correctness of the answer. The same set of criteria is used for manual annotation.
Question Literacy
Does the question exhibit correct grammar, spelling, and punctuation? This criterion assesses the linguistic quality of the question. A well-formed question should be free of typographical errors, contain appropriate punctuation, and follow standard grammatical rules. Additionally, the phrasing should align grammatically with the surrounding context, ensuring the question does not feel syntactically out of place.
Question Clarity
Is the intent of the question clear and unambiguous? This criterion evaluates how easily a reader can understand what information is being requested. The question should be interpretable either based on the provided context or general knowledge, without requiring additional clarification. Vague, overly broad, or logically inconsistent questions should be penalized.
Question Naturalness
Does the question sound like it could have been written by a human? This assesses whether the question appears natural and contextually appropriate. It should avoid signs of being artificially generated such as unnatural phrasing, rigid templates, or repetitive structures. A natural question should feel relevant and plausible within the discourse of the text.
Context Sufficiency
Can the answer to this question be found entirely within the provided context? This criterion determines whether the context passage contains enough information to answer the question. A question should not require external knowledge or assumptions unless that knowledge is very general or trivial. Questions with answers that are clearly present and verifiable in the text should receive high marks.
Context Necessity
Is the provided context necessary to answer the question? This evaluates whether the question meaningfully engages with the context. Ideal questions should be context-dependent, meaning they cannot be accurately answered without access to the specific passage. Generic or overly broad questions that could be answered independently of the text (without specialized knowledge) are discouraged.
Answer Literacy
Is the answer written in a grammatically correct and readable manner? This criterion checks for the overall linguistic quality of the answer. It should be free from spelling mistakes, awkward constructions, or inconsistent grammatical structure.
Answer Correctness
Is the answer factually correct and appropriate for the given question? This criterion gauges the accuracy of the generated answer. It should contain all possible entities that can be mentioned in the answer without omitting any necessary details.
Answer Uniqueness Based on Context
Is this the only plausible answer that can be given based on the text? This checks whether the answer is uniquely determined by the information in the context. If the passage contains multiple plausible answers or if ambiguity remains, this checkbox should not be selected. Ideal answers should be both correct and exclusive given the text.
Appendix G LLM-as-Judge filtering
For fine-grained filtering of the generated question-answer pairs, we employ the LLM-as-a-Judge paradigm, leveraging a POLLUX 7B Martynov et al. (2025) model to systematically evaluate each sample along a diverse set of criteria. This approach offers a scalable and cost-effective alternative to human annotation , while still maintaining a high level of judgment quality.
Every criterion from Appx. F is transformed into a separate prompt with the specific scoring scale (0-2). For answer criteria (Literacy, Correctness, Uniqueness Based on Context), the prompt contains a news article, a question, and an answer; for other criteria, the answer is omitted. The example is classified as positive according to the particular criterion if the judge model assigns a rating of 1 or higher. This threshold was chosen to imitate the majority vote used in human evaluation.
To validate the reliability of using a language model as an automated evaluator for filtering generated question-answer pairs, we conducted an empirical comparison against human judgments. A random sample of 532 examples was drawn from the generated dataset and independently assessed by a panel of human annotators (with more than three annotators per example) as well as by a large language model. An example was considered positive by human annotators if half or more of the assessors provided a positive assessment.
| Criterion | Precision | Recall |
| Question Literacy | 0.96 | 0.99 |
| Question Clarity | 0.99 | 0.62 |
| Question Naturalness | 0.96 | 0.52 |
| Context Sufficiency | 0.94 | 0.71 |
| Context Necessity | 0.93 | 0.95 |
| Answer Correctness | 0.95 | 0.82 |
| Answer Uniqueness Based on Context | 0.85 | 0.78 |
| Answer Literacy | 0.91 | 0.97 |
The comparison in Tab. 6 reveals that the language model achieves high Precision but moderate Recall relative to the human-labeled data. This trade-off is acceptable in our setting, as the dataset contains a large volume of generated examples. In this context, precision is more critical than recall: retaining only high-quality samples is preferable, even if some potentially acceptable data are discarded. This justifies the use of the language model as an effective filter for selecting the most reliable and contextually appropriate question-answer pairs at scale.
Appendix H Experimental Setup Details
H.1 Embedding Model Prefixes
To vectorize questions and documents, we used embedders with the corresponding prefixes. These prefixes are shown in Tab. 7.
| Model | Query prefix | Text prefix |
| FRIDA | search_query: | search_document: |
| E5 Mistral | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
| Qwen 3 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
| mE5 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
H.2 LLM Prompt Template
To generate answers for the questions, we used the following template for the user message prompt:
‘‘Answer the question using the provided context.
Give me only an answer.
<context> {context} </context>
Question: {question}
Answer: ’’
H.3 Model Configuration
For serving models, we used the vLLM framework. The used model parameters are shown in the Tab. 8.
| Model | tp_size | max_context_length |
| Qwen 2.5 | 4 | 32768 |
| Qwen 2.5 | 1 | 32768 |
| Ruadapt Qwen | 4 | 32768 |
| Zero Mistral | 4 | 131072 |
| Gemma 3 | 1 | 131072 |
| Gemma 3 | 4 | 131072 |
We set max_new_tokens to 1000 for all models to limit the response length of the models.
H.4 Metric Description
The performance of retrieval is measured by 3 metrics:
Hit Rate measures the proportion of queries for which relevant document appears among the top-k retrieved results.
Recall computes the ratio of relevant documents successfully retrieved out of all ground-truth relevant documents.
Normalized Discounted Cumulative Gain (NDCG) evaluates the ranking quality by assigning higher scores to relevant documents retrieved at higher ranks.
We evaluate End-to-end RAG systems with:
ROUGE-L, used to assess generation quality, is based on the longest common subsequence between the model output and reference text.
Substring Matching measures whether key segments from the reference answer are present in the generated response.
Judge Score is used to evaluate the overall answer quality, is calculated as the average of the automatic scores from Pollux 151515https://ai-forever.github.io/POLLUX/ across multiple criteria (e.g., correctness, completeness, and relevance).
Appendix I Human Evaluation Interface
A screenshot of a system used for human evaluation is presented in Fig. 6.
Appendix J LLM-as-Judge RAG Evaluation
This section provides a detailed description of the LLM-as-Judge criteria used to evaluate RAG systems. To build a comprehensive and interpretable set of metrics, Evaluation Targets provided by Yu et al. (2024) are utilized. For each Evaluation Target, we select several criteria from the POLLUX Martynov et al. (2025) set of criteria:
Answer Relevance. Measures the alignment between the generated response and the content of the initial query.
Absence of unnecessary details. (Fluff) The LLM’s output is relevant and do not contain fluff.
Faithfulness. Estimates the quality of the information extraction from retrieved documents.
Consistency with real-world facts. The LLM’s output does not contain factual errors.
Correctness of results. The LLM extracted correct information from the text.
Correctness Measures the accuracy of the generated response by comparing it to the ground truth response.
Completeness. The answer is complete and reaches the goal.
Factual accuracy. The LLM correctly reproduced the necessary facts and their related context.
Preserving the main idea and details of the original. The LLM preserves details and main idea.
The fine-grained set of metrics allows for comparing the RAG systems more precisely and improves interpretability. Fig. 6 provides a comparison of different RAG systems built on the basis of different variants of the Qwen 2.5 model combined with FRIDA and Qwen 3 Embedding 8B retrieval models. The results clearly demonstrate that larger language models yield higher-quality responses across all criteria, while the Absence of unnecessary details criterion results are similar for all combinations. Additionally, systems using Qwen3 8B embeddings consistently outperform those using FRIDA, highlighting the critical role of retrieval quality in end-to-end RAG performance. These findings emphasize that both the generative and retrieval components contribute significantly to final system effectiveness.

### v2

@@
 By design, it can potentially be adapted to other languages and multilingual settings, broadening its applicability beyond Russian.
-(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
+(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
 2 Related Work
@@
 (2024) in the RAG evaluation landscape such as temporal aspects of information, dynamic and vast knowledge sources, and factuality and faithfulness in generation.
-Architecture of the Dynamic Benchmark
+Dynamic Benchmark Architecture
 The entire pipeline of the benchmark architecture is shown in Fig. 2. Below, each step is described in more detail.
-Data Acquisition and Processing: We have maintained a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx.
+Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx.
 A for the news sources list) on a daily basis.
@@
 4.
-The pipeline transforms generated QA pairs into several HF datasets, which form the core of our benchmark:
+The pipeline transforms the generated QA pairs into several HF datasets, which form the core of our benchmark:
 Public Texts: Contains cleaned source documents. Each item is assigned a public_id to enable matching without exposing the true internal IDs.
@@
 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation.
-The KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph.
+KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph.
 The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
@@
 Vrandečić and Krötzsch (2014).
-To perform this, we search by API in this open knowledge base for the entities extracted in the previous step.
+To do this, we search by API in this open knowledge base for the entities extracted in the previous step.
 Afterward, the found entities are vectorized using an embedding model and saved as a vector database.
@@
 2 demonstrate consistently strong performance across all evaluated retriever models.
-Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par except for MRR, where E5 Mistral7b Instruct beats its competitor.
-However, mE5lInstruct and FRIDA also perform competitively.
+Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par.
+mE5lInstruct also performs competitively.
+As for FRIDA, it also demonstrates strong performance but its results are slightly inferior to those of the competitors.
 | Retriever | Hit Rate | Recall | NDCG |
@@
 We set max_new_tokens to 1000 for all models to limit the response length of the models.
-H.4 Metric Description
+H.4 Metrics
 The performance of retrieval is measured by 3 metrics:
@@
 A screenshot of a system used for human evaluation is presented in Fig. 6.
-Appendix J LLM-as-Judge RAG Evaluation
+Appendix J LLM-as-Judge RAG Evaluation Criteria
 This section provides a detailed description of the LLM-as-Judge criteria used to evaluate RAG systems.
@@
 (2024) are utilized.
-For each Evaluation Target, we select several criteria from the POLLUX Martynov et al.
-(2025) set of criteria:
+For each Evaluation Target, we select several criteria from the POLLUX set of criteria:
 Answer Relevance. Measures the alignment between the generated response and the content of the initial query.

### v3

@@
-DRAGON: Dynamic RAG Benchmark On News
+DRAGOn: Designing RAG On Periodically Updated Corpus
 Abstract
-Retrieval-Augmented Generation (RAG) is a widely adopted approach for improving the factuality of large language models (LLMs) by incorporating external knowledge at inference time.
-Although there exist multiple RAG benchmarks for English, evaluation resources for other languages, including Russian, remain scarce and static, failing to capture the dynamic nature of real-world deployments.
-In this work, we present DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating RAG systems in Russian on a changing news corpora.
-DRAGON is built upon a regularly updated corpus of Russian news and public documents and supports comprehensive evaluation of both the retriever and generator components.
-Question generation is performed automatically with the use of Knowledge Graph constructed from the corpus and enables the extraction of four core question types aligned with distinct subgraph patterns.
-We release a complete evaluation framework comprising the pipeline for automatic question generation, evaluation scripts, which are potentially reusable for other languages and multilingual settings, and benchmark data.
-We also launch a public leaderboard to encourage community participation and comparison.
-DRAGON: Dynamic RAG Benchmark On News
+This paper introduces DRAGOn, method to design a RAG benchmark on a regularly updated corpus.
+It features recent reference datasets, a question generation framework, an automatic evaluation pipeline, and a public leaderboard.
+Specified reference datasets allow for uniform comparison of RAG systems, while newly generated dataset versions mitigate data leakage and ensure that all models are evaluated on unseen, comparable data.
+The pipeline for automatic question generation extracts the Knowledge Graph from the text corpus and produces multiple question-answer pairs utilizing modern LLM capabilities.
+A set of diverse LLM-as-Judge metrics is provided for a comprehensive model evaluation.
+We used Russian news outlets to form the datasets and demonstrate our methodology.
+We launch a public leaderboard to track the development of RAG systems and encourage community participation.
+DRAGOn: Designing RAG On Periodically Updated Corpus
+Fedor Chernogorskii2,1, Sergei Averkiev1, Liliya Kudraleeva3, Zaven Martirosian1,4,8, Maria Tikhonova1,5, Valentin Malykh6,3,7, Alena Fenogenova1,5 1SberAI, 2MBZUAI, 3ITMO, 4MISIS, 5HSE University, 6MWS AI, 7IITU, 8YSDA Correspondence: fechernogor@gmail.com
 1 Introduction
-Retrieval-Augmented Generation (RAG) has become a powerful instrument for enhancing the domain adaptation and factuality of large language models (LLMs) by incorporating external knowledge retrieved at inference time.
+Retrieval-Augmented Generation (RAG) has become a powerful tool for enhancing the domain adaptation and factuality of large language models (LLMs) by incorporating external knowledge retrieved at inference time.
 This approach enables more up-to-date and grounded responses without the need for costly re-training.
-As RAG-based systems expand to applications such as open-domain QA, customer support, and enterprise search, their standardized evaluation becomes crucial for assessing reliability and utility.
-While several RAG benchmarks have been proposed for the English language, there are significantly fewer evaluation resources for other languages, including Russian.
-In this work, we focus on the evaluation of RAG systems for the Russian language.
-We introduce DRAGON: Dynamic RAG Benchmark On News (Fig.
-1) a novel benchmark that reflects realistic usage patterns by leveraging a regularly updated knowledge base derived from current news sources.
+As RAG-based systems expand to applications such as open-domain QA, customer support, and enterprise search, their standardized evaluation remains a challenge.
+It may be unclear whether strong performance of a system is due to the quality of its retriever-generator pipeline or because the underlying LLM has been exposed to portions of the test data during training.
+It is possible that a static benchmark will become contaminated over time.
+Several existing RAG evaluation frameworks Es et al.
+(2024); Lyu et al.
+(2025) provide pipelines for automatic generation of question-answer pairs, typically assuming that users will build their own evaluation datasets.
+While this enables domain-specific benchmarking, it is labor-intensive and difficult to maintain.
+Furthermore, when the retrieval corpus is continuously updated during deployment, results may become non-reproducible, as the model may face a different knowledge distribution than the one used during its initial evaluation.
+In this work, we introduce DRAGOn: Designing RAG On Periodically Updated Corpus a novel methodology that reflects realistic usage patterns by leveraging a regularly updated knowledge base.
+We also release a benchmark built on current news sources using this methodology, which can be readily adapted to other document domains, such as scientific papers or court decisions.
 To foster transparency and community engagement, we publicly release an evaluation framework which comprises the codebase for automatic question generation, evaluation scripts, and a dynamic leaderboard to track progress on RAG-based systems in Russian.
-While the benchmark targets Russian, the framework is potentially extendable to other languages and multilingual scenarios, making it broadly applicable.
+Although the benchmark targets Russian, the framework is potentially extendable to other languages and multilingual scenarios, making it broadly applicable.
 Our contributions are as follows:
-(i) We propose DRAGON111The video demonstration of the evaluation tool is available on YouTube., the first RAG benchmark with a regularly updated knowledge base based on news corpora for Russian, designed to evaluate RAG systems in a dynamic setup.
+(i) We propose DRAGOn111The video demonstration of the evaluation tool is available on YouTube.
+the methodology to develop a RAG benchmark with a regularly updated knowledge base, designed to evaluate RAG systems in a dynamic setup; we develop a benchmark on Russian news corpora as a reference for the proposed methodology.
 (ii) We release an open-source evaluation framework 222The framework is released under the MIT license: https://github.com/RussianNLP/DRAGON comprising a reusable question generation pipeline and evaluation scripts, enabling reproducible experimentation and easy integration of new models and retrieval components.
 By design, it can potentially be adapted to other languages and multilingual settings, broadening its applicability beyond Russian.
-(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
+(iii) We launch a regularly updated public leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
 2 Related Work
-Proper assessment of RAG systems requires holistic benchmarks that evaluate both the retrieval and the generation components.
-For English, a growing body of research has focused on designing such benchmarks across a range of reasoning and grounding challenges.
-KILT Petroni et al.
-(2021) provides a unified benchmark over a fixed Wikipedia snapshot, with an emphasis on source attribution and retrieval grounding.
-Conversational settings with multi-turn dialogue scenarios are investigated in mtRAG Katsis et al.
+Evaluating retrieval-augmented generation (RAG) systems poses unique challenges, as it requires datasets that jointly assess both the retrieval and generation components.
+Constructing such benchmarks is costly and time-consuming because it involves curating large collections of text–question–answer triplets.
+To alleviate this, several works have explored synthetic data generation to automate question and answer creation Es et al.
+(2024); Lyu et al.
+(2025), often leveraging domain-specific pipelines or knowledge graphs for better control over content and difficulty.
+Early RAG benchmarks such as KILT Petroni et al.
+(2021) unified multiple English-language datasets over a fixed Wikipedia snapshot, emphasizing source attribution and retrieval grounding.
+More recent efforts have extended the evaluation to multi-turn conversational and reasoning-intensive scenarios, as seen in mtRAG Katsis et al.
 (2025) and RAD-Bench Kuo et al.
 (2025).
-The factual consistency of generated responses in RAG has become a critical challenge.
-CRAG Yang et al.
-(2024) introduces a comprehensive benchmark based on five main RAG characteristics.
-RAGAS Es et al.
-(2024) proposes a reference-free evaluation suite that captures context relevance, faithfulness, and answer completeness, along with an open-source API for practical benchmarking.
-Despite these advances, most of the RAG benchmarks focus on English.
-Evaluation for other languages remains limited.
-NoMIRACL Thakur et al.
-(2024) provides multilingual QA tasks in 18 languages and analyzes retrieval failures and hallucination rates.
-For Russian, RusBEIR Kovalev et al.
-(2025) offers retrieval-only evaluation, but a complete end-to-end RAG benchmark has yet to be developed.
-To address this need we present DRAGON – a dynamic, regularly updated benchmark for Russian-language RAG systems based on real-world, shifting corpora.
-3 System Demo
-The benchmark is designed to evaluate RAG systems in a dynamically evolving news domain.
-Its architecture prioritizes modularity, automation, and reproducibility while addressing the core challenges Yu et al.
-(2024) in the RAG evaluation landscape such as temporal aspects of information, dynamic and vast knowledge sources, and factuality and faithfulness in generation.
-Dynamic Benchmark Architecture
-The entire pipeline of the benchmark architecture is shown in Fig. 2. Below, each step is described in more detail.
-Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx.
-A for the news sources list) on a daily basis.
-Newly parsed content is synchronized into our storage.
+The CRAG benchmark Yang et al.
+(2024) further focuses on factual consistency, capturing five key aspects of RAG system behavior.
+Complementarily, RAGAS Es et al.
+(2024) provides a reference-free evaluation framework measuring context relevance, faithfulness, and answer completeness, and offers an open-source API for reproducible benchmarking.
+Dynamic and time-sensitive evaluation has emerged as another important dimension.
+RealTime QA Kasai and et al.
+(2023) introduced a benchmark for evaluating systems on continuously evolving information sources, reflecting real-world deployment settings.
+The news domain, with its frequent updates and temporal drift, has thus become a popular testbed for such studies Tang and Yang (2024); Chen et al.
+(2024).
+Despite these advances, the field still lacks a universal, contamination-free, and continuously updated benchmarks White et al.
+(2024).
+This gap hinders fair and reproducible comparison across RAG systems and motivates the development of dynamic, standardized evaluation resources.
+To address this need, we present DRAGOn – a methodology to develop dynamic, regularly updated benchmarks for RAG systems based on real-world, shifting corpora.
+3 Benchmark Design
+Using our methodology, one can develop a benchmark designed to evaluate RAG systems in a dynamically evolving news domain.
+The benchmark’s architecture prioritizes modularity, automation, and reproducibility while addressing the core challenges Yu et al.
+(2024) in the RAG evaluation landscape, such as the temporal aspects of information, the vast and dynamic sources of knowledge, and the factuality and faithfulness in generation.
+The entire pipeline of the benchmark architecture is shown in Fig.
+2.
+Below, each step is described in more detail.
+Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia on a daily basis.
+The parsed content is synchronized with our storage.
 To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
-This design ensures the benchmark reflects evolving real-world distributions and mitigates the risks of overfitting to static datasets.
+This design ensures that the benchmark reflects evolving real-world distributions and mitigates the risks of overfitting to static datasets.
 The pipeline further ensures that newly surfaced topics and entities from the news stream are constantly incorporated into the benchmark.
-QA Dataset Formation: The process of creating the questions and answers based on the updated increment of news data is described in detail in Sec.
+QA Dataset Formation: The process of creating questions and answers based on the updated increment of the news data is described in detail in Sec.
 4.
-The pipeline transforms the generated QA pairs into several HF datasets, which form the core of our benchmark:
+The pipeline transforms the generated QA pairs into several HF datasets, which form the core of the benchmark:
 Public Texts: Contains cleaned source documents. Each item is assigned a public_id to enable matching without exposing the true internal IDs.
@@
 This enables convenient local experimentation, debugging, and reproducibility.
-All datasets are versioned and uploaded to Hugging Face with incrementally updated revision numbers.
-This versioning mechanism ensures reproducibility and provides users with stable snapshots for further experimentation.
 User Experience
-To facilitate seamless evaluation for users, we provide a PyPi-hosted Python library rag_bench, which offers an interface to:
-Fetch the latest version of the public datasets by dynamically resolving the latest Hugging Face revision; Observe the RAG system baseline which can be adopted for the target one; Evaluate RAG system and package results for submission; Submit results via API to our evaluation portal; Calculate retrieval and generative metrics locally using the sandbox datasets.
+To facilitate seamless evaluation for users, we provide a PyPi-hosted Python library rag_bench, which offers an interface to: Fetch the latest version of the public datasets by dynamically resolving the latest Hugging Face revision; Observe the RAG system baseline, which can be adopted for the target one; Evaluate RAG system and package results for submission; Submit results via API to our evaluation portal; Calculate retrieval and generative metrics locally using the sandbox datasets.
 User workflow includes loading public data, applying a custom RAG pipeline, and collecting results in the following form:
@@
 "found_ids": [17, 69, 69, 22, ...],
-"model_answer": "Answer": "Top"
+"model_answer": "Answer"
 These results encode both the retrieved public_ids and the generated answers, decoupling the user’s model output from any private evaluation artifacts.
-This separation allows secure evaluation without exposing ground-truth data.
+This separation allows for secure evaluation without exposing ground-truth data.
 Validation Portal
@@
 Leaderboard and Auto-Evaluation
-A Hugging Face Gradio Space serves as the public Leaderboard (see Appx.
-B for a screenshot).
-Results are committed in a version-controlled results.json file, automatically updated by the validation portal upon approval.
-To reduce latency and improve benchmarking coverage, we support automatic evaluation for selected pre-approved baselines, which include several popular LLMs and retrieval embedding models (see Appx.
-C for the details).
+A Hugging Face Gradio Space serves as the public Leaderboard. The results are committed in a version-controlled results.json file, automatically updated by the validation portal upon approval.
+To reduce latency and improve benchmarking coverage, we support automatic evaluation for selected pre-approved baselines, which include several popular LLMs and retrieval embedding models.
 The results are computed via the same rag_bench client.
@@
 Note that sandbox datasets are not updated on a regular basis. They serve as a static reference set for demonstration and local validation purposes.
-4 Methodology
-The Data Generation pipeline (Fig.
+4 Dataset Generation
+The Data Generation pipeline (see Fig.
 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation.
 KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph.
-The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
-Knowledge Graph Extraction
-To enable greater control over question generation, we developed a Knowledge Graph Extraction module, inspired by Chepurova et al. (2024).
-At the first stage, the LLaMa 3.3 70B Instruct444https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Grattafiori et al.
-(2024) is used for the extraction of triplet-candidates.
-Each triplet consists of a head entity or subject, a relation, and a tail entity or object.
-All the extracted entities are used to match with entities from Wikidata555We use Russian subgraph of Wikidata as a base one.
-Vrandečić and Krötzsch (2014).
-To do this, we search by API in this open knowledge base for the entities extracted in the previous step.
-Afterward, the found entities are vectorized using an embedding model and saved as a vector database.
-Then, for each entity extracted in the first step, we list five of the closest entity candidates in the vector representation base.
-At the final stage, we use the same LLaMa 3.3 70B Instruct for normalization and unification of the extracted entities and relations.
-For each entity and relation in the triplet, the language model analyzes the list of candidates obtained at the previous stage and decides on the possibility of replacing the original entity or relation with one of the candidates.
-The triplet context is taken into account by adding an original text.
-This approach allows the unification of naming variants for the same entity, which is critical for the subsequent construction of the knowledge graph.
-We aim to create a graph that contains novel information.
-To ensure this, we discard any triplets that are completely matched with saved entities and relations from the open knowledge base (Wikidata).
-We assume that facts not included in Wikidata are novel and thus unlikely to have been seen by the language model.
-Question Generation
-The question generation stage begins with subgraph extraction.
-We systematically sample all subgraphs that match one of the four predefined structural templates, each corresponding to one of the question types.
-We employ a typology of four question types, inspired by Yang et al.
-(2024): simple, set, multi-hop, and conditional questions (see Appx.
-D for a detailed description).
-A subgraph is passed to the language model, which is instructed to generate both a natural language question and a corresponding answer.
-The question is formulated as a coherent sentence, while the answer consists of one or more entities derived from the subgraph.
-QA Filtering
-The dataset requires filtering to improve the quality of question-answer pairs and mitigate language model errors.
-Firstly, the linguistic acceptability of each question is measured using a model trained on the RuCoLa dataset666https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola Mikhailov et al.
-(2022) model, which scores sentences based on their grammatical correctness and fluency.
-Secondly, we extract Named Entities from the original document, solving the NER task with Natasha library.
-We look for these entities in generated question-answer pairs.
-The pairs that do not contain named entities are discarded to avoid examples relying on general knowledge.
-The questions are passed to the small LLMs (namely, we use Qwen 2.5 7B Instruct Team (2024), Llama 3 8B Grattafiori et al.
-(2024)) without context to detect the simplistic pairs that do not challenge the language model.
-Thirdly, each generated pair is checked for compliance with the source subgraph.
-Either question or answer should contain an entity from the subgraph.
-To ensure this, we search the closest by Levenshtein distance Levenshtein et al.
-(1966) entity from the subgraph.
-We consider an entity to be found if the smallest distance from the entities in the subgraph is under the predefined threshold.
-We assume that the generated set should contain all the entities from the subgraph and not contain any extra ones.
-Otherwise, there is a generation error or the question type mismatch (see Appx.
-E).
-The final stage uses the LLM-as-Judge approach to measure the question and answer quality.
-The pretrained judge POLLUX 7B Martynov et al.
-(2025) specifically trained for a fine-grained evaluation in Russian is adopted for this task.
-The model evaluates each sample against 8 generative criteria developed to assess the quality of the question-answer pair and its context dependence (the set description is given in Appx.
-F).
-As a result, questions of low quality are excluded from the final test set.
-The evaluation process is described in Appx.
-G.
-After the filtering stage 150 questions for each category form the final test dataset.
+The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with an LLM.
+4.1 Knowledge Graph Extraction
+To achieve fine-grained control over automatic question generation, we designed a Knowledge Graph (KG) Extraction module inspired by Chepurova et al.
+(2024).
+This component transforms unstructured news texts into a structured set of factual triplets that later guide question creation.
+We use LLaMa 3.3 70B Instruct444https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Grattafiori et al.
+(2024) to extract candidate factual triplets from the corpus.
+Each triplet has the form (head entity – relation – tail entity), corresponding to the subject, predicate, and object in the original sentence.
+The extracted entities are matched with the Russian subgraph of Wikidata Vrandečić and Krötzsch (2014).
+For every entity name identified in the text, we query the Wikidata API to find possible matches.
+The mapped entities are then vectorized using a sentence-embedding model and stored in a vector database.
+To handle ambiguity, we keep the five most similar candidates for each extracted entity according to vector similarity.
+To ensure consistency across triplets, the same LLaMa 3.3 70B Instruct model is used again to normalize entity and relation names.
+Given the list of candidate matches from the previous step, the model selects the most appropriate canonical form while taking the full sentence context into account.
+This process merges spelling variants and aliases referring to the same entity, resulting in a cleaner, unified graph structure.
+Our goal is to build a graph that captures new information appearing in the latest news updates.
+Therefore, we discard any triplets that exactly match existing facts in Wikidata.
+Triplets absent from the knowledge base are treated as novel facts, as they are more likely to represent fresh events, and thus serve as valuable material for generating time-sensitive questions.
+4.2 Question Types
+The question generation stage begins with subgraph extraction from the constructed Knowledge Graph.
+We identify all subgraphs matching one of four predefined structural templates, each representing a distinct question type Yang et al.
+(2024):
+Simple
+These type correspond the most simple questions based a single fact mentioned in one or several texts.
+They are based only on one relation from the graph: the predicate and one of the entities involved in the relation are used to compose the question, and the second entity becomes the answer.
+relations: (Morty Smith | voice | Keisuke Chiba)
+question: Who voiced Morty Smith?
+answer: Keisuke Chiba
+Set
+Set questions test the RAG system’s ability to align information from several texts.
+They are based on a one-to-many subgraphs in which the number of triplets share relation and either object or subject.
+The question is generated using shared entity and relation.
+The answer consists of all other entities in the subgraph.
+relations: (Ryan Otter | composed music for | Method), (Ryan Otter | composed music for | Trigger)
+question: What projects has Ryan Otter composed music for?
+answer: Trigger, Method
+Multi-Hop
+Multi-hop questions evaluate the system’s ability to reason in a multistage manner.
+The corresponding subgraph is a pair of triplets, intersecting at a single entity.
+The question is constructed similarly to a simple question; however, the repeated entity must not be mentioned in the question.
+It is used as a bridge-entity, which is described in question as a reference extracted from another triplet.
+relations: (FAW | country of origin | China), (FAW | number of cars sold in 2023 | 2139)
+question: In which country is the company located that sold 2139 cars in 2023?
+answer: China
+Conditional
+Conditional questions are the extension of multi-hop questions with the same underlying subgraph of a pair of triplets, intersecting at a single entity.
+However, for a conditional question, both facts are used to form the question, while the repeated entity becomes an answer.
+relations: (Roman Miroshnichenko | performed at | M-bar), (Roman Miroshnichenko | met with | Dmitry Dibrov)
+question: Who performed at M-bar and met with Dmitry Dibrov?
+answer: Roman Miroshnichenko
+Each selected subgraph is then passed to the language model, which generates a natural-language question–answer pair.
+The question is formulated as a fluent, contextually appropriate sentence, while the answer comprises one or more entities explicitly present in the subgraph.
+4.3 QA Filtering
+To ensure high-quality and contextually grounded question–answer pairs, we apply a multi-stage filtering pipeline combining linguistic validation, entity consistency, graph correspondence check, and LLM-based judgment.
+1. Linguistic and Structural Filtering.
+Firstly, we assess the grammatical correctness and fluency of each question using a RuRoBERTa-large model trained on the RuCoLa dataset555https://huggingface.co/RussianNLP/ruRoBERTa-large-rucola Mikhailov et al.
+(2022).
+This step eliminates ungrammatical or poorly formed questions.
+Next, we perform Named Entity Recognition (NER) on the original source text using the Natasha library.
+The generated questions and answers are checked for the presence of these entities.
+Samples without explicit named entities are discarded to remove trivial, knowledge-free examples.
+We further filter out overly simplistic questions by evaluating them with smaller instruction-tuned LLMs (Qwen 2.5 7B Team (2024) and LLaMa 3 8B Grattafiori et al.
+(2024)) without context; if a model can answer a question directly from prior knowledge, it is excluded.
+2. Graph Correspondence Filtering.
+Each remaining QA pair is verified against the source subgraph used during question generation to ensure factual alignment.
+For every entity in the subgraph, we calculate the Levenshtein distance Levenshtein and others (1966) between its label and the text of both the question and answer.
+Each node in the graph (entity) is assigned 2 coefficients: question presence and answer presence.
+It is evaluated as the scaled Levenshtein distance between the name of the entity and the closest substring from the question and answer.
+These values allow us to check that all entities have been mentioned correctly.
+In the graphs for Set and Conditional question types, the positions of every entity are strictly determined.
+The algorithm averages the presence coefficients of entities implied to be in the same part of the output.
+If any of these values is lower than the threshold, it indicates incorrect generation.
+For Simple questions each entity can appear in both parts of the output, although the entity must be mentioned once in the question-answer pair.
+The presence coefficients were averaged over all entities from the subgraph, then 5% highest and lowest values were filtered out.
+Multi-Hop questions inherit the same process for nodes having only one connection in the subgraph.
+The bridge entity that has two connections should not be mentioned in the model output.
+A high value for any of the presence coefficients for this entity demonstrates a question-type violation.
+3. LLM-as-Judge Evaluation.
+In the final stage, we apply the LLM-as-Judge approach using POLLUX 7B Martynov et al.
+(2025), a model fine-tuned for fine-grained evaluation in Russian.
+Each QA pair is automatically rated along eight generative criteria: (1) Question literacy (grammar and style), (2) Clarity, (3) Naturalness, (4) Context sufficiency (answer can be found in the passage), (5) Context necessity (question depends on the passage), (6) Answer correctness, (7) Answer uniqueness, and (8) Answer literacy.
+More details on these criteria can be found in Appx.
+D.
+Each criterion is transformed into a separate prompt with the specific scoring scale (0-2).
+For answer criteria (Literacy, Correctness, Uniqueness Based on Context), the prompt contains a news article, a question, and an answer; for other criteria, the answer is omitted.
+The example is classified as positive according to the particular criterion if the judge model assigns a rating of 1 or higher.
+This threshold was chosen to imitate the majority vote used in human evaluation.
+To validate the reliability of using a language model as an automated evaluator for filtering generated question-answer pairs, we conducted an empirical comparison against human judgments.
+A random sample of 532 examples was drawn from the generated dataset and independently assessed by a panel of human annotators (with more than three annotators per example) as well as by a large language model.
+An example was considered positive by human annotators if half or more of the assessors provided a positive assessment.
+| Criterion | Precision | Recall |
+| Question Literacy | 0.96 | 0.99 |
+| Question Clarity | 0.99 | 0.62 |
+| Question Naturalness | 0.96 | 0.52 |
+| Context Sufficiency | 0.94 | 0.71 |
+| Context Necessity | 0.93 | 0.95 |
+| Answer Correctness | 0.95 | 0.82 |
+| Answer Uniqueness | 0.85 | 0.78 |
+| based on context | ||
+| Answer Literacy | 0.91 | 0.97 |
+The comparison in Tab.
+1 reveals that the language model achieves high Precision but moderate Recall relative to the human-labeled data.
+This trade-off is acceptable in our setting, as the dataset contains a large volume of generated examples.
+In this context, precision is more critical than recall: retaining only high-quality samples is preferable, even if some potentially acceptable data are discarded.
+This justifies the use of the language model as an effective filter for selecting the most reliable and contextually appropriate question-answer pairs at scale.
+After all filtering stages, 150 high-quality questions per category are retained for the final benchmark dataset.
 5 Experimental Setup
-To construct our experimental RAG systems, we used the LangChain framework777https://pypi.org/project/langchain/.
+To construct our experimental RAG systems, we used the LangChain framework666https://pypi.org/project/langchain/.
 All texts from the Public Texts dataset are split into chunks of length 500 with an overlap of 100 characters.
-Each chunk is vectorized using the retrieval model of the evaluating RAG system with the corresponding document prefixes (see Appx.
-H.1 for their exact formulation), and the resulting vectors are stored in a vector database.
+Each chunk is vectorized using the retrieval model of the evaluating RAG system with the corresponding document prefixes, and the resulting vectors are stored in a vector database.
 During the search phase, we use the prompted retrieval model to find five of the most relevant texts that match the user’s query.
-Retrieved chunks are incorporated into a prompt (see Appx.
-H.2) provided to the LLM of the evaluated RAG system.
+Retrieved chunks are incorporated into a prompt provided to the LLM of the evaluated RAG system.
 If the total length of the filled-in prompt exceeds the model’s maximum context length, the contextual information is truncated to the required size.
-To accelerate LLM inference, we utilize the vLLM framework888https://github.com/vllm-project/vllm Kwon et al.
+To accelerate LLM inference, we utilize the vLLM framework777https://github.com/vllm-project/vllm Kwon et al.
 (2023).
-Other configuration details are described in Appx.
-H.3.
-Scoring metrics
-The performance of retrieval is measured with the 3 main metrics: Hit Rate, Recall, and NDCG.
-End-to-end RAG-system evaluation is performed via ROUGE-L, Substring Matching (SM), and Judge Score.
-See Appx.
-H.4 for their description.
+5.1 Experimental Setup Details
+Embedding Model Prefixes
+To vectorize questions and documents, we used embedders with the corresponding prefixes. These prefixes are shown in Tab. 2.
+| Model | Query prefix | Text prefix |
+| FRIDA | search_query: | search_document: |
+| E5 Mistral | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
+| Qwen 3 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
+| mE5 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
+LLM Prompt Template
+To generate answers for the questions, we used the following template for the user message prompt:
+‘‘Answer the question using the provided context.
+Give me only an answer.
+<context> {context} </context>
+Question: {question}
+Answer: ’’
+Model Configuration
+For serving models, we used the vLLM framework. The model parameters used are shown in Tab. 3.
+| Model | TP | ML |
+| Qwen 2.5 | 4 | 32768 |
+| Qwen 2.5 | 1 | 32768 |
+| Ruadapt Qwen | 4 | 32768 |
+| Qwen 3 | 4 | 32768 |
+| Gemma 3 | 1 | 131072 |
+| Gemma 3 | 4 | 131072 |
+We set max_new_tokens to 1000 for all models to limit the response length of the models.
+Metrics
+The performance of retrieval is measured by 3 metrics:
+Hit Rate measures the proportion of queries for which the relevant document appears among the top-k retrieved results.
+Mean Reciprocal Rank (MRR) evaluates ranking quality by measuring how highly the first relevant document is ranked, assigning higher scores when a relevant document appears earlier in the ranked list.
+We evaluate End-to-end RAG systems with:
+ROUGE-2 measures bigram overlap between the model output and the reference text, capturing local phrase-level similarity and rewarding matching adjacent word pairs.
+ROUGE-L, measures the longest common subsequence between the model output and the reference text, capturing overlap in overall sentence structure.
+The Judge Score is used to evaluate the overall answer quality, is calculated as the average of the automatic scores from Pollux 888https://ai-forever.github.io/POLLUX/ across multiple criteria (e.g., correctness, completeness, and relevance).
 6 Experiments
@@
 Each QA pair from Sandbox Datasets (Sec.
-2) is independently evaluated by 3 expert annotators along the evaluation criteria from Appx.
-F.
-Annotators were asked to mark each pair as “Good” or “Not Good” with respect to each criterion (an example of the assessment instruction is described in Appx.
-I).
+2) is independently evaluated by 3 expert annotators along the evaluation criteria from Sec.
+D.
+Annotators were asked to mark each pair as “Good” or “Not Good” with respect to each criterion.
 To account for potential subjectivity in judgment, we considered a QA pair to be acceptable with the majority vote.
-Tab.
-1 shows the proportion of QA pairs considered good for each dataset version and each evaluation criterion.
-The results establish the high quality of generated questions and significant context dependency.
-The answer evaluation proved the prevalence of correct answers, while the answer uniqueness is lower, so the ground truth answer can be substituted with another entity from the text.
-This fact exhibits the importance of LLM-as-Judge evaluation for RAG systems to avoid rephrasing influence.
-| Criterion | Apr | May | Jun | |
-| Question | Question Literacy | 0.96 | 0.97 | 0.99 |
-| Clarity | 0.99 | 1.00 | 1.00 | |
-| Naturalness | 0.98 | 0.96 | 0.97 | |
-| Context Sufficiency | 0.98 | 0.98 | 0.99 | |
-| Context Necessity | 0.95 | 0.97 | 0.98 | |
-| Answer | Correctness | 0.95 | 0.92 | 0.96 |
-| Uniqueness | 0.76 | 0.78 | 0.80 | |
+Evaluation results are provided in Tab.
+4.
+| Criterion | Apr | May | Jun |
+| Question Literacy | 0.96 | 0.97 | 0.99 |
+| Clarity | 0.99 | 1.00 | 1.00 |
+| Naturalness | 0.98 | 0.96 | 0.97 |
+| Context Sufficiency | 0.98 | 0.98 | 0.99 |
+| Context Necessity | 0.95 | 0.97 | 0.98 |
+| Correctness | 0.95 | 0.92 | 0.96 |
+| Uniqueness | 0.76 | 0.78 | 0.80 |
 | Answer Literacy | 0.79 | 0.71 | 0.75 |
@@
 5 demonstrate consistently strong performance across all evaluated retriever models.
-Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par.
-mE5lInstruct also performs competitively.
+Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores.
+mE5Large Instruct also performs competitively.
 As for FRIDA, it also demonstrates strong performance but its results are slightly inferior to those of the competitors.
-| Retriever | Hit Rate | Recall | NDCG |
-| FRIDA | 0.88 | 0.84 | 0.85 |
-| mE5 | 0.90 | 0.87 | 0.86 |
-| Qwen 3 | 0.93 | 0.90 | 0.89 |
-| E5 Mistral | 0.93 | 0.89 | 0.87 |
+| Retriever | Hit Rate | MRR |
+| FRIDA | 0.932 | 0.822 |
+| Qwen 3 | 0.960 | 0.867 |
+| E5 Mistral | 0.956 | 0.851 |
+| mE5 | 0.949 | 0.834 |
 End-to-End System Evaluation
-Tab. 4 presents general results of the full RAG-system evaluation and detailed criteria-wise evaluation scores are given in Appx. J.
-First, it can be seen that the choice of the retrieval model plays a crucial role and same as in Sec.
-6.
+| LLM | Rouge2 | RougeL | JS |
+| Retrieval: FRIDA | |||
+| Gemma 3 12B it | 0.14 | 0.22 | 0.63 |
+| Gemma 3 27B it | 0.14 | 0.22 | 0.64 |
+| Qwen 2.5 32B | 0.09 | 0.16 | 0.62 |
+| Qwen 2.5 7B | 0.08 | 0.13 | 0.57 |
+| Qwen3 32B | 0.08 | 0.14 | 0.64 |
+| Rudadapt Qwen 32B | 0.13 | 0.22 | 0.72 |
+| Retrieval: mE5 | |||
+| Gemma 3 12B it | 0.15 | 0.24 | 0.67 |
+| Gemma 3 27B it | 0.15 | 0.24 | 0.67 |
+| Qwen 2.5 32B | 0.10 | 0.18 | 0.66 |
+| Qwen 2.5 7B | 0.09 | 0.15 | 0.63 |
+| Qwen3 32B | 0.11 | 0.18 | 0.69 |
+| Rudadapt Qwen 32B | 0.14 | 0.21 | 0.74 |
+| Retrieval: Qwen 3 | |||
+| Gemma 3 12B it | 0.16 | 0.26 | 0.71 |
+| Gemma 3 27B it | 0.17 | 0.26 | 0.72 |
+| Qwen 2.5 32B | 0.11 | 0.19 | 0.68 |
+| Qwen 2.5 7B | 0.09 | 0.16 | 0.64 |
+| Qwen3 32B | 0.11 | 0.19 | 0.71 |
+| Rudadapt Qwen 32B | 0.16 | 0.25 | 0.82 |
+| Retrieval: E5 Mistral | |||
+| Gemma 3 12B it | 0.16 | 0.25 | 0.68 |
+| Gemma 3 27B it | 0.16 | 0.25 | 0.70 |
+| Qwen 2.5 32B | 0.12 | 0.19 | 0.68 |
+| Qwen 2.5 7B | 0.09 | 0.16 | 0.64 |
+| Qwen3 32B | 0.11 | 0.19 | 0.71 |
+| Rudadapt Qwen 32B | 0.16 | 0.25 | 0.79 |
+The results are provided in Table 6. Overall, the results show that classic metrics such as Rouge-L are not objective enough and do not allow evaluating all aspects of the RAG task.
+First, it can be seen that the choice of the retrieval model plays a crucial role.
 Qwen3 and E5 Mistral show the strongest results.
-Second, it should be noted that the general LLM ranking remains the same with every retrieval, with Qwen 2.5 and Zero Mistral heading the list by Judge Score and SM, and Gemma 3 over-performing other competitors by Rouge-L.
-Overall, the results show that classic metrics such as Rouge-L are not objective enough and do not allow evaluating all aspects of the RAG task.
+Second, it should be noted that the general LLM ranking remains the same with every retrieval, with Rudadapt Qwen 32B heading the list by Judge Score, and Gemma 3 outperforming other competitors by Rouge metrics.
 In general, system scores positively characterize DRAGON as being complex enough for modern RAG-systems, allowing researchers to evaluate their capabilities at a high level.
 In the future, we also plan to complexify Judge Evaluation criteria, thus providing an opportunity for an adequate assessment of more advanced models than those that exist nowadays and avoiding the danger of the benchmark being solved.
-| Retriever | LLM | Rouge-L | SM | Judge Score |
-| FRIDA | Qwen 2.5 | 0.42 | 0.47 | 0.83 |
-| Qwen 2.5 | 0.33 | 0.43 | 0.80 | |
-| Ruadapt Qwen | 0.40 | 0.45 | 0.82 | |
-| Zero Mistral | 0.44 | 0.45 | 0.83 | |
-| Gemma 3 | 0.44 | 0.44 | 0.81 | |
-| Gemma 3 | 0.40 | 0.47 | 0.81 | |
-| Qwen 3 | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
-| Qwen 2.5 | 0.36 | 0.45 | 0.83 | |
-| Ruadapt Qwen | 0.42 | 0.48 | 0.86 | |
-| Zero Mistral | 0.45 | 0.48 | 0.86 | |
-| Gemma 3 | 0.47 | 0.47 | 0.84 | |
-| Gemma 3 | 0.43 | 0.49 | 0.85 | |
-| E5 Mistral | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
-| Qwen 2.5 | 0.36 | 0.46 | 0.82 | |
-| Ruadapt Qwen | 0.41 | 0.47 | 0.85 | |
-| Zero Mistral | 0.45 | 0.47 | 0.85 | |
-| Gemma 3 | 0.46 | 0.47 | 0.83 | |
-| Gemma 3 | 0.41 | 0.48 | 0.85 | |
-| mE5 | Qwen 2.5 | 0.43 | 0.49 | 0.85 |
-| Qwen 2.5 | 0.34 | 0.44 | 0.82 | |
-| Ruadapt Qwen | 0.39 | 0.46 | 0.83 | |
-| Zero Mistral | 0.44 | 0.46 | 0.84 | |
-| Gemma 3 | 0.46 | 0.46 | 0.83 | |
-| Gemma 3 | 0.41 | 0.48 | 0.83 |
 7 Conclusion
-We presented DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating retrieval-augmented generation systems in Russian.
-DRAGON reflects real-world deployment settings by leveraging a regularly updated knowledge base and focuses on the recurrent evaluation of both retriever and generator components.
-Our benchmark addresses the current lack of standardized RAG evaluation tools for the Russian language.
-We release the benchmark, the framework comprising a question generation pipeline and evaluation scripts, and launch a public leaderboard, to support reproducible, transparent, and community-driven research.
+We presented DRAGOn, a method to design RAG benchmark on any periodically updated document source, with it we created the dynamic benchmark for evaluating retrieval-augmented generation systems in Russian.
+DRAGOn is designed for real-world deployment settings by leveraging a regularly updated knowledge base and focusing on the recurrent evaluation of both retriever and generator components.
+Our methodology addresses the current lack of standardized RAG evaluation tools; thus, we created a sample benchmark for the Russian language.
+We release the benchmark, which comprises a question generation pipeline and evaluation scripts, and launch a public leaderboard to support reproducible, transparent, and community-driven research.
 In the future, with the evolving capabilities of RAG systems, we plan to extend the benchmark by introducing new question types, refining the LLM-as-Judge criteria.
@@
 The proposed benchmark consists entirely of Russian language documents and questions.
-Although the methodology itself could be easily applied to any other language, in the current state the only one language is presented.
+Although the methodology itself could be easily applied to any other language, in its current state, only one language is presented.
 Evaluation Metrics
 The chosen evaluation metrics, such as ROUGE, which is essentially an n-gram precision, predominantly focus on surface-level matching.
-These metrics may not adequately reflect the semantic and pragmatic aspects of the generated content, and have limited correlation with human judgment Deutsch et al.
+These metrics may not adequately reflect the semantic and pragmatic aspects of the generated content and have limited correlation with human judgment Deutsch et al.
 (2022).
-LLM as Judge evaluation is designed to mitigate a semantic gap of n-gram based metrics.
-However, RAG benchmark requires specific criteria to catch system performance details.
+The LLM as Judge evaluation is designed to mitigate the semantic gap of n-gram-based metrics.
+However, the RAG benchmark requires specific criteria to capture details of system performance.
 Building more adapted judge models can improve the quality of the assessment.
@@
 Future work will continue to refine these frameworks to address emerging ethical challenges as the field evolves.
+Error Analysis
+A further limitation of our current benchmark release is the lack of a systematic error analysis of model failures.
+While we report aggregate retrieval and generation scores, we do not yet provide a fine-grained breakdown of common failure modes (e.g., retrieval misses vs.
+ranking issues, incomplete evidence aggregation in multi-hop questions, hallucinations under partially relevant context).
+Such analysis is important both for interpreting leaderboard progress and for understanding whether improvements come from better retrieval, better grounding/faithfulness, or exploiting dataset artifacts.
+In future work, we plan to add structured error taxonomies and identify systematic weaknesses of evaluated RAG pipelines.
 AI-assistants Help
@@
 Nevertheless, some segments of our publication can be potentially detected as AI-generated, AI-edited, or human-AI-generated.
+Acknowledgments
+We would like to express our deep appreciation to Ivan Bondarenko for his contribution to our pipeline and generous support of this work.
+This research partially done by A.F.
+is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University).
 Appendix A News Data Sources
@@
 | Ruadapt Qwen | 32B | msu-rcc-lair/RuadaptQwen-32B-instruct | Tikhomirov and Chernyshev (2024) |
-| Zero Mistral | 24B | ZeroAgency/Zero Mistral-24B | – |
-| Gemma 3 | 12B | google/gemma-3-12b-it | Team (2025) |
-| Gemma 3 | 27B | google/gemma-3-27b-it | Team (2025) |
+| Qwen 3 | 32B | Qwen/Qwen3-32B | Team (2025b) |
+| Gemma 3 | 12B | google/gemma-3-12b-it | Team (2025a) |
+| Gemma 3 | 27B | google/gemma-3-27b-it | Team (2025a) |
 | Model | Size | Hugging Face Hub link | Citation |
@@
 | mE5 | 560M | intfloat/multilingual-e5-large-instruct | Wang et al. (2024) |
-Appendix D Question Types
-We use the typology of four question types inspired by Yang et al. (2024). Below each type is described in more detail:
-Simple
-These type correspond the most simple questions based a single fact mentioned in one or several texts.
-They are based only on one relation from the graph: the predicate and one of the entities involved in the relation are used to compose the question, and the second entity becomes the answer.
-relations: (Morty Smith | voice | Keisuke Chiba)
-question: Who voiced Morty Smith?
-answer: Keisuke Chiba
-Set
-Set questions test the RAG system’s ability to align information from several texts.
-They are based on a one-to-many subgraphs in which the number of triplets share relation and either object or subject.
-The question is generated using shared entity and relation.
-The answer consists of all other entities in the subgraph.
-relations: (Ryan Otter | composed music for | Method), (Ryan Otter | composed music for | Trigger)
-question: What projects has Ryan Otter composed music for?
-answer: Trigger, Method
-Multi-Hop
-Multi-hop questions evaluate the system’s ability to reason in a multistage manner.
-The corresponding subgraph is a pair of triplets, intersecting at a single entity.
-The question is constructed similarly to a simple question; however, the repeated entity must not be mentioned in the question.
-It is used as a bridge-entity, which is described in question as a reference extracted from another triplet.
-relations: (FAW | country of origin | China), (FAW | number of cars sold in 2023 | 2139)
-question: In which country is the company located that sold 2139 cars in 2023?
-answer: China
-Conditional
-Conditional questions are the extension of multi-hop questions with the same underlying subgraph of a pair of triplets, intersecting at a single entity.
-However, for a conditional question, both facts are used to form the question, while the repeated entity becomes an answer.
-relations: (Roman Miroshnichenko | performed at | M-bar), (Roman Miroshnichenko | met with | Dmitry Dibrov)
-question: Who performed at M-bar and met with Dmitry Dibrov?
-answer: Roman Miroshnichenko
-Appendix E Graph correspondence filtering
-The question-answer pairs were compared to the graph used for generation.
-Each node in the graph (entity) is assigned 2 coefficients: question presence and answer presence.
-It is evaluated as the scaled Levenshtein distance between the name of the entity and the closest substring from the question and answer.
-These values allow us to check that all entities have been mentioned correctly.
-In the graphs for Set and Conditional question types, the positions of every entity are strictly determined.
-The algorithm averages the presence coefficients of entities implied to be in the same part of the output.
-If any of these values is lower than the threshold, it indicates incorrect generation.
-For Simple questions each entity can appear in both parts of the output, although the entity must be mentioned once in the question-answer pair.
-The presence coefficients were averaged over all entities from the subgraph, then 5% highest and lowest values were filtered out.
-Multi-Hop questions inherit the same process for nodes having only one connection in the subgraph.
-The bridge entity that has two connections should not be mentioned in the model output.
-A high value for any of the presence coefficients for this entity demonstrates a question-type violation.
-Appendix F Question-Answer Evaluation Criteria
+Appendix D Question-Answer Evaluation Criteria
 This section describes the question-answer evaluation criteria used on the final question filtering stage.
@@
 Ideal answers should be both correct and exclusive given the text.
-Appendix G LLM-as-Judge filtering
-For fine-grained filtering of the generated question-answer pairs, we employ the LLM-as-a-Judge paradigm, leveraging a POLLUX 7B Martynov et al.
-(2025) model to systematically evaluate each sample along a diverse set of criteria.
-This approach offers a scalable and cost-effective alternative to human annotation , while still maintaining a high level of judgment quality.
-Every criterion from Appx.
-F is transformed into a separate prompt with the specific scoring scale (0-2).
-For answer criteria (Literacy, Correctness, Uniqueness Based on Context), the prompt contains a news article, a question, and an answer; for other criteria, the answer is omitted.
-The example is classified as positive according to the particular criterion if the judge model assigns a rating of 1 or higher.
-This threshold was chosen to imitate the majority vote used in human evaluation.
-To validate the reliability of using a language model as an automated evaluator for filtering generated question-answer pairs, we conducted an empirical comparison against human judgments.
-A random sample of 532 examples was drawn from the generated dataset and independently assessed by a panel of human annotators (with more than three annotators per example) as well as by a large language model.
-An example was considered positive by human annotators if half or more of the assessors provided a positive assessment.
-| Criterion | Precision | Recall |
-| Question Literacy | 0.96 | 0.99 |
-| Question Clarity | 0.99 | 0.62 |
-| Question Naturalness | 0.96 | 0.52 |
-| Context Sufficiency | 0.94 | 0.71 |
-| Context Necessity | 0.93 | 0.95 |
-| Answer Correctness | 0.95 | 0.82 |
-| Answer Uniqueness Based on Context | 0.85 | 0.78 |
-| Answer Literacy | 0.91 | 0.97 |
-The comparison in Tab.
-6 reveals that the language model achieves high Precision but moderate Recall relative to the human-labeled data.
-This trade-off is acceptable in our setting, as the dataset contains a large volume of generated examples.
-In this context, precision is more critical than recall: retaining only high-quality samples is preferable, even if some potentially acceptable data are discarded.
-This justifies the use of the language model as an effective filter for selecting the most reliable and contextually appropriate question-answer pairs at scale.
-Appendix H Experimental Setup Details
-H.1 Embedding Model Prefixes
-To vectorize questions and documents, we used embedders with the corresponding prefixes. These prefixes are shown in Tab. 7.
-| Model | Query prefix | Text prefix |
-| FRIDA | search_query: | search_document: |
-| E5 Mistral | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
-| Qwen 3 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
-| mE5 | Instruct: Given a web search query, retrieve relevant passages that answer the query. Query: | X |
-H.2 LLM Prompt Template
-To generate answers for the questions, we used the following template for the user message prompt:
-‘‘Answer the question using the provided context.
-Give me only an answer.
-<context> {context} </context>
-Question: {question}
-Answer: ’’
-H.3 Model Configuration
-For serving models, we used the vLLM framework. The used model parameters are shown in the Tab. 8.
-| Model | tp_size | max_context_length |
-| Qwen 2.5 | 4 | 32768 |
-| Qwen 2.5 | 1 | 32768 |
-| Ruadapt Qwen | 4 | 32768 |
-| Zero Mistral | 4 | 131072 |
-| Gemma 3 | 1 | 131072 |
-| Gemma 3 | 4 | 131072 |
-We set max_new_tokens to 1000 for all models to limit the response length of the models.
-H.4 Metrics
-The performance of retrieval is measured by 3 metrics:
-Hit Rate measures the proportion of queries for which relevant document appears among the top-k retrieved results.
-Recall computes the ratio of relevant documents successfully retrieved out of all ground-truth relevant documents.
-Normalized Discounted Cumulative Gain (NDCG) evaluates the ranking quality by assigning higher scores to relevant documents retrieved at higher ranks.
-We evaluate End-to-end RAG systems with:
-ROUGE-L, used to assess generation quality, is based on the longest common subsequence between the model output and reference text.
-Substring Matching measures whether key segments from the reference answer are present in the generated response.
-Judge Score is used to evaluate the overall answer quality, is calculated as the average of the automatic scores from Pollux 151515https://ai-forever.github.io/POLLUX/ across multiple criteria (e.g., correctness, completeness, and relevance).
-Appendix I Human Evaluation Interface
+Appendix E Human Evaluation Interface
 A screenshot of a system used for human evaluation is presented in Fig. 6.
-Appendix J LLM-as-Judge RAG Evaluation Criteria
+Appendix F LLM-as-Judge RAG Evaluation Criteria
 This section provides a detailed description of the LLM-as-Judge criteria used to evaluate RAG systems.